OpenStudy (shes.radical):

I really do not get this :/ A bucket of paint has spilled on a tile floor. The paint flow can be expressed with the function r(t) = 3t, where t represents time in minutes and r represents how far the paint is spreading. The flowing paint is creating a circular pattern on the tile. The area of the pattern can be expressed as A(r) = πr2. Part A: Find the area of the circle of spilled paint as a function of time, or A[r(t)]. Show your work. Part B: How large is the area of spilled paint after 10 minutes? You may use 3.14 to approximate π in this problem.

Nnesha (nnesha):

A[r(t)] means substitute r for r(t) function into A(r)

OpenStudy (shes.radical):

...what? so then its \[A[r(t)]=3.14(10^2)+3t ?\] i dont get it..

Nnesha (nnesha):

don't add r(t) = what ?

OpenStudy (shes.radical):

3t

Nnesha (nnesha):

yes right so you can replace r(t) with 3t A(3t) =pi r^2 replace r with 3t don't add 3t

OpenStudy (shes.radical):

oh..is that it?

Nnesha (nnesha):

yes right \[\huge\rm A(\color{red}{r(t)})=A(\color{ReD}{3t})= \pi\color{Red}{ r}^2\] ^replace r with 3t and then take square

OpenStudy (shes.radical):

....so then its \[A(3t)=3.14*3^2?\]

Nnesha (nnesha):

not just 3 it's (3t)^2

OpenStudy (shes.radical):

okay so then its \[A(3t)=3.14(3t^2)\] now what .-.

Nnesha (nnesha):

(3t)^2 means 3^2 * t^2

OpenStudy (shes.radical):

which is \[9(t^2)\]

Nnesha (nnesha):

yes right \[\huge\rm A(\color{red}{r(t)})=A(\color{ReD}{3t})= 9t^2\pi\]

Nnesha (nnesha):

part B substitute t for 10

Nnesha (nnesha):

use 3.14 for pi

OpenStudy (shes.radical):

\[9(10^2)(3.14)=2826\]

Nnesha (nnesha):

looks good

OpenStudy (shes.radical):

so, Part A: \[A[r(t)]=A(3t)=9t^2 \pi\] and Part B: \[9(10^2)(3.14)=2826\] right? thats the answer right?? :D

Nnesha (nnesha):

yep

Nnesha (nnesha):

or you can say 9pit^2 :D

OpenStudy (shes.radical):

you're awesome..thank you thank you thank you!!

Nnesha (nnesha):

np :=)